### v1

# Can Charge Transfer Across C–H···O Hydrogen Bonds Stabilize Oil Droplets in Water?

Abstract
Oil-water emulsions resist aggregation due to the presence of negative charges at their surface that leads to mutual repulsion between droplets, but the molecular origin of oil charge is currently under debate. While much evidence has suggested that hydroxide ions must accumulate at the interface, an alternative perspective attributes the negative charge on the oil droplet to not an ionic species, but charge transfer of electron density from water to oil molecules. While the charge transfer mechanism is consistent with the correct sign of oil charge based on electrophoresis experiments, it is just as important to provide good estimates of the magnitude of the negative charge to explain emulsion stability. Here we show using energy decomposition analysis that the amount of net flow of charge from water to oil is negligibly small due to nearly equal forward and backward charge transfer through weak oil-water interactions, such that oil droplets would be unstable and coalesce, contrary to experiment. The lack of charge transfer also explains why vibrational sum frequency scattering reports a blue shift in the oil C-H frequency when forming emulsions with water, which arises from Pauli repulsion due to localized confinement at the interface.
1Kenneth S. Pitzer Theory Center and Department of Chemistry, 2Departments of Bioengineering and 3Chemical and Biomolecular Engineering, University of California, Berkeley, Berkeley, CA, 94720 USA
4Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA, 94720 USA
corresponding author: thg@berkeley.edu
1 Introduction
Oil and water are capable of forming stable emulsions in which the minor component forms droplets dispersed throughout the solvent. Oil-water emulsions are thought to be electrostatically stabilized by surface charges that lead to mutual repulsion between droplets to confer stability. For the case of oil droplets in water, electrophoresis experiments consistently show that the oil droplets migrate toward the anode under an applied electric field1, 2, which has typically been interpreted to mean that the oil droplets preferentially adsorb negative ions present in the aqueous solution. This gives rise to a stationary layer of ions that adhere to the oil particle and travel with the droplet under the applied field, creating an electrostatic potential difference with the mobile fluid at the “slip plane” which defines the $\zeta$ potential. The $\zeta$ potential can be inferred from the electrophoretic mobility of the droplet and is frequently used as a gauge of emulsion stability3, 4, with $\zeta$ potential magnitudes less than $\sim$30 mV resulting in flocculation or coalescence5. For oil-in-water emulsions prepared at neutral pH, the $\zeta$ potential is typically around --60 to --90 mV,6, 7, 8 indicating excellent long term droplet stability.
Because stable oil-in-water emulsions can be prepared without added ions, it is frequently assumed that hydroxide ions from the auto-dissociation of water preferentially adsorb to the water oil-interface.9, 10 Indeed, the $\zeta$ potential varies strongly with the pH, including going through an isoelectric point around pH = 3.1 In 2004, Beattie and Djerdjev estimated the surface charge density of interfacial ions by combining electroacoustic droplet sizing with measurements of the NaOH required to maintain a constant pH during homogenization.11 They observed that the amount of OH- added increased linearly with increasing interfacial area, suggesting an intrinsic and constant surface charge density at the oil-water interface. Under the assumption of excess hydroxide ions they estimated the surface charge to be --0.28 e-/nm2. More recently using second harmonic generation (SHG) experiments, Gan et al. found that hydroxide, not hydronium, showed preferential absorption at the surface of an hexadecane emulsion in water.10 Their estimates of hydroxide surface charge of --0.14 e-/nm2, which is comparable to the value obtained previously by Beattie and Djerdjev, would support an adsorption free energy of $\sim$-∼ -8.3 kcal/mol.10
However a different perspective has been offered over the last 15 years that attributes the negative charge on the oil droplet not to ionic species, but an accumulation of excess electron density at the interface arising from charge transfer (CT).6 In 2011 Vacha and co-workers6 combined mobility measurements, sum frequency scattering (SFS) experiments, and molecular dynamics simulations. The mobility measurements measured a $\zeta$ potential of --81 mV, while the SFS experiments showed no selective adsorption of hydroxide ions to the oil-water interface. The molecular dynamics (MD) simulations found that a small amount of charge was transferred between water hydrogen bonds, not between water and oil, using the TIP4P-FQ water model that is capable of modeling CT.12 They calculated a water surface charge density of $-0.0001$ to $-0.0028$ e-/nm2, and subsequent studies employing similar charge equilibration models13, 14 also reported comparable surface charge densities in the range of $-0.001$ to $-0.002$ e-/nm2. However, the water surface charge estimates discussed in these studies primarily refer to CT interactions between water molecules rather than CT between water and oil.6, 13, 14.
In subsequent work Pullanchery and co-workers introduced vibrational sum frequency scattering (VSFS) measurements on oil-water emulsions15, and found spectroscopic signatures in which the oil C–H modes blue shift while the water vibrations red shift at the interface. This spectroscopic signature was claimed to be supportive for a CT mechanism, since they again found no experimental evidence from VSFS of interfacial OH radicals or OH- species.15 Supporting theoretical calculations further proposed that CT occurs across C–H···O hydrogen bonds15. The CT mechanism has also been analyzed from theoretical calculations using the density-derived electrostatic and chemical (DDEC)16 atomic charge partitioning scheme, which found a charge of --0.015 e-/nm2, but only after accumulation into the droplet and deep in the oil phase.17 This is an order of magnitude higher than previous CT calculations.6, 13, 14 More recently in 2024, Roke and co-workers found that while oil mobility changes as a function of increasing pH, the blue-shifted vibrational signatures of the C–H peak remain unperturbed by increasing pH which they claim confirms a CT mechanism18. To explain the mobility results, they conducted simulations and found that under a 0.1 V/ Å electric field oriented along the z-axis, neopentane exhibited a net drift relative to the surrounding water which they attributed not solely to CT but also to field-induced ’dynamic polarization’, to help explain the sensitivity of mobility, but not vibrations, to pH.18
While the CT offers an intriguing explanation for interfacial charging of the oil phase, there are several reasons to doubt its viability for explaining stability, electrokinetic experiments, and spectroscopic measurements of oil-water emulsions. To address these concerns, we employ model cluster calculations as well as a polarizable force field simulation of an extended water–hexane interface to quantify the amount of CT using quantum mechanical energy decomposition analysis based on the absolutely localized molecular orbital (ALMO-EDA) method19, 20, 21, 22, 23. Our findings indicate that the surface charge densities computed by ALMO-EDA are one to two orders of magnitude smaller than the DDEC atomic charge partitioning scheme or charges obtained from Mulliken populations. Furthermore, true CT analysis shows any CT from water to oil is largely negated by back-transfer of electron density24 from oil to water, in which back-transfer is a natural energy lowering process when one correctly accounts for electron flow. Therefore such small amounts of net CT can’t possibly explain oil-water emulsion stability and hence does not explain oil charging and electrophoresis experiments. Finally, we reaffirm again that vibrational blue shifts of C-H actually arise from Pauli repulsion whereas CT would redshift these same modes25, 26, 22, and thus a CT mechanism can’t be the origin of the VSFS spectroscopic observations. Finally we show that explanations based on dynamic polarization to explain mobility18 lack evidence due to simulation errors that fail to exhibit directional drift under an applied electric field.
2 Results
Beginning with accurate state-of-the-art density functional theory (DFT) calculations27, the ALMO-EDA framework19, 20, 21, 22, 23 decomposes the overall intermolecular interaction energy, $\Delta E_{\text{INT}}$ between several species or fragments into three physically interpretable contributions, providing insight into their relative significance.
$\Delta E_{\text{INT}}=\Delta E_{\text{FRZ}}+\Delta E_{\text{POL}}+\Delta E_{\text{CT}}$ (1)
[…]
Within the ALMO-EDA framework, electron density rearrangement associated with CT can be interpreted as donor-acceptor orbital interactions. CT occurs when an occupied orbital on one molecule donates electron density to a complementary virtual orbital on another molecule. This interaction is quantified through complementary occupied-virtual pair (COVP)19, 28, 29 analysis. In what follows we consider a single dominant forward COVP and a single dominant backward COVP, because they are sufficient to capture approximately 75% of the overall CT contribution to both the energy and net charge, making them a compact summary of the CT orbital picture. A more detailed description of the ALMO-EDA method, as implemented in Q-Chem30, and further details of forward and backward CT using COVP analysis are provided in the Methods section.
2.1 Charge transfer across the C-H···O hydrogen bond
We begin with the prototype system for understanding CT based on the strong hydrogen bond O-H···O formed by the the water dimer, in order to compare it to C-H···O interactions in which it is claimed that CT occurs from water to oil15. As seen in Figure 1A, the COVP indicates that hydrogen bonding in the water dimer is composed of CT from an electron pair localized on an oxygen atom donating electrons to the O–H $\sigma$ anti-bonding orbital on the other water. For the water dimer, CT occurs predominantly in the forward direction, from the hydrogen-bond acceptor water molecule to the donor water molecule, with a net CT of 2.67 me- quantified in full by ALMO-EDA.
To accurately assess CT in oil–water interactions, we next analyze the C-H···O interaction that has been widely recognized as another, albeit weaker, type of hydrogen bond.25, 31 Many studies impose a linear constraint when analyzing C-H···O interactions25, however their relatively weak strength makes them more susceptible to bending and less sensitive to deviations from their equilibrium bond length25. This is seen in the optimized structures of a hexane molecular fragment (C6H14) interacting with one water molecule as shown in Figure 1B, in which the optimized structure exhibits a C-H···O bond angle of approximately 120°, with the hydrogens in water positioned closer to hexane than the oxygen atoms. This significantly deviates from the linear hydrogen bond configuration observed in the water dimer or idealized C-H···O configurations.
The accompanying dominant COVPs from ALMO-EDA show that forward CT (0.25 me-) from the $\pi$ type lone pair electrons of water’s oxygen is an order of magnitude smaller than for the water-water hydrogen bond because the C-H group is a poor donor for hydrogen-bonding. Moreover, a backward CT of 0.26 me- from hexane to the $\sigma^{\star}$ orbital of water nearly compensates for the forward CT. Given the close spatial proximity of water hydrogen atoms to hexane, electron density flow from the oxygen of water toward hexane while simultaneously being drawn back from hexane to the hydrogen atom of water, results in a negligible 0.01 me- transfer of charge. Even if we are to include all forward and backward COVP components, Figure 1D indicates that the total net CT only amounts to 0.10 me-, i.e, 25 times smaller than the water dimer. Consequently, it is not surprising that the optimized hexane-water structure does not exhibit the typical linear hydrogen-bonding configuration of the water dimer that is strengthened by a large net CT interaction.
Figure 1: Charge transfer analysis for optimized water–hexane complexes compared to the water dimer. Optimized structures and dominant COVPs in forward and backward charge transfer processes of A (H2O)(H2O), B (C6H14)(H2O), and C(C6H14)(H2O1—H2O2) complexes. Forward CT refers to H2O $\rightarrow$ C6H14, and backward CT refers to C6H14 $\rightarrow$ H2O. D and E represent charge transfer and ALMO-EDA energy analysis of optimized complexes in A, B, C, and C1 and C2 that represent the complexes (C6H14)(H2O1) and (C6H14)(H2O2) after removing one water molecule from complex C. All structural optimizations were performed at the $\omega$B97X-V/def2-TZVP level, and AMLO-EDA calculations were conducted using $\omega$B97X-V/def2-TZVPD. CT unit me-. The original data for D and E are provided in Supplementary Tables S1 and S2.
In Figure 1C we see that the presence of a second water molecule enhances the forward CT by a factor of $\sim$1.8 compared to the hexane-water dimer in Figure 1B. This enhancement arises from the formation of a stronger electron donor through hydrogen bonding with the second water molecule. However, the increased forward CT is accompanied by an increase in backward CT, as the second water molecule becomes more electron-deficient, with the virtual orbitals in the backward CT primarily localized on the second water molecule. In this case the C-H group returns electron density to the anti-bonding orbital of the second water given the close range of H··H interactions. To further investigate the role of the C-H···O hydrogen bond geometry in the CT mechanism, we remove one water molecule from the (C6H14) (H2O1—H2O2) to create (C6H14)(H2O1) (C1) and removing the other water molecule to create (C6H14)(H2O2) (C2). As seen in Figure 1C, C1 has a well-defined hydrogen bond with a C··O distance of $\sim$3.35Å, resulting in a net CT of 0.36 me-, whereas C2 lacks an optimal hydrogen bond configuration, leading to a net CT in the reverse direction from hexane to water of --0.28 me-, which explains the negligible net charge flow in Figure 1C. These findings indicate that even though the presence of an additional water molecule enhances forward CT, the overall net electron transfer remains negligibly small, and furthermore it accumulates on water not hexane.
Finally it is important to emphasize that both forward and backward CT processes stabilize the system and contribute substantially to lowering its energy. As shown in Figure 1E, the CT interaction energy plays a similarly crucial role in stabilizing the two water–hexane complex as it does in the water dimer. But unlike the water dimer, forward and backward CT density are nearly equal in magnitude but opposite in sign, resulting in a mostly balanced charge exchange between water and hexane. This challenges the assumption that the presence of a C-H···O hydrogen bond necessarily induces significant, directional charge transfer from water to oil.
2.2 Charge transfer at an extended oil-water interface
Analysis of minimized dimer and trimer configurations reveal that forward and backward CT between water and hexane nearly cancels. However, the CT mechanism at the extended water–oil interface may be significantly more intricate due to thermal fluctuations and a more extended network of hydrogen-bond interactions between waters and hydrophobic association of hexanes. To gain further insight into the CT process at an extended water–oil interface, molecular dynamics at ambient conditions using the polarizable AMOEBA force field32, 33 was employed; simulation details are provided in the Methods section.
From the MD simulations we draw two types of interfacial hexane-water clusters. Figure 2A is a representative snapshot of the 200 clusters in which the water molecules are localized around multiple hexanes, whereas Figure 2B is a representative snapshot of the 50 clusters in which interfacial water molecules form more extensive hydrogen-bonded network around the hexane. From these clusters we derive the probability density distribution in Figure 2C that shows a high-probability region near a CH…O angle of 120° for C-O distances $>$ 3.5 Å, consistent with the C-H···O configurations of the ab initio optimized clusters in Figure 1B,C. The broad distribution means that the C-H···O hydrogen bond is significantly weaker and more distorted compared to the highly directional hydrogen bonds found for even thermalized water–water interactions which experience significant CT. This would indicate that CT in the forward direction from water to oil will be much smaller at an extended oil-water interface as well. In Figure 2D, the distribution of distances between hexane and water hydrogens is centered at 1.85 Å, whereas the distribution of water oxygen to hexane hydrogen distances is peaked at 2.2 Å, and the two distributions exhibit minimal overlap. This interfacial organization implies that the hydrogen of surface water orientated towards the oil (“dangling O-H groups”) will be a conduit for back-transfer of electron density from oil to water.
This is borne out in Figure 2E which illustrates the ensemble nature of CT dynamics at the oil-water interface. The forward CT distribution of the interfacial clusters peaks at 2.09 me-/nm2, which arises from the C-H···O hydrogen bond at the interface, and similar to the small clusters in Figure 1. But again this is balanced by electron backflow from the CT distribution centered around -1.44 me-/nm2 due to the presence of short-range H···H
Figure 2: Water/oil interface structures. (A, B) Sample representation of the clusters considered for the ALMO-EDA calculations highlighting the spatial arrangement of water molecules (red and white spheres) around hexane molecules (cyan). The extended hydrogen-bonded water network at the interface is illustrated in pink. (C) Probability density distribution illustrating the spatial and angular relationship between water oxygen and hexane. The x-axis represents the C–O distance, while the y-axis shows the C-H···O angle, describing the relative orientation of water molecules with respect to hexane’s C–H bonds. (D) Probability distribution of the closest distance between hydrogen atoms in hexane with water hydrogens and oxygens. This is consistent with the radial distribution functions (Supplementary Figure S1), where the H···H intermolecular distribution is more short-ranged than the O···H distribution in the water–hexane system. (E) Probability density distributions of Forward CT, Backward CT, and Net CT. Solid lines for direct water–hexane interactions clusters in (A) and dashed lines correspond to the clusters in (B). The original data is provided in Supplementary Table S3 and S4. (F) Probability density distributions of water oil inter-molecule interaction energy components. The legend includes the percentage contribution of each energy component relative to the total interaction energy. Forward CT refers to H2O $\to$ C6H14, and Backward CT refers to C6H14 $\to$ H2O. Cluster selection details are provided in the Methods section.
interactions at the interface that defines the source of backward CT as also seen in Figure 1. The average net CT of 0.00065 e-/nm2 (0.65 me-/nm2) obtained from ALMO-EDA corresponds to 0.00007 e-/hydrogen bond, closely matching the 0.0001 e-/hydrogen bond observed in the hexane–water clusters in Figure 1.
Furthermore these conclusions are consistent between the clusters with localized waters near hexane or the clusters incorporating the extended hydrogen bonding network since interfacial CT is predominantly dictated by direct water–hexane interactions, whereas the broader hydrogen bonding network within the interfacial region exerts only a subtle perturbative influence (average net CT of 0.00061 e-/nm2). Finally Figure 2F shows that the frozen energy term constitutes the dominant component (61.6%) to water-hexane interactions, with similar percentages of polarization energy and CT making up the remaining energetic interactions. The estimated 20.3% of CT energy between water and hexane is a proportion similar to that observed in Figure 1 for small water–hexane clusters B and C, and is energetically the manifestation of electron delocalization in stabilizing the system.
2.3 Surface charge and oil-water emulsion stability
Using the non-linear Poisson Boltzmann equation34, 35, 36, 37, we consider how different estimates of surface charge give rise to changes in surface potential which is a measure of emulsion stability. The relationship between the surface potential $\psi_{0}$ and the surface charge density $\sigma_{0}$ can be expressed as:
$\sigma_{0}=Z\sqrt{2\varepsilon_{r}\varepsilon_{0}k_{B}T\rho_{\text{bulk}}}\left[2\sinh\left(\frac{e\psi_{0}}{2k_{B}T}\right)+\frac{4}{R\left(\frac{\varepsilon_{0}\varepsilon_{r}k_{B}T}{2e^{2}Z^{2}\rho_{\text{bulk}}}\right)^{-1/2}}\tanh\left(\frac{e\psi_{0}}{4k_{B}T}\right)\right],$ (2)
where $\varepsilon_{r}$ is the relative permittivity, $k_{B}$ is the Boltzmann constant, $T$ is the temperature, $e$ is the elementary charge, and $R$ is the droplet radius. We note that previous experiments have reported average surface droplet radii of 100 nm15, 8 to 230 nm38, 18; here we use a radius of 125 nm as reported most recently18, 8. We also consider the relationship between surface charge and surface potential under different ionic concentrations, where Z is the ion valence and $\rho_{\text{bulk}}$ is the bulk ion density. While a concentration of hydroxide ions of 0.0001 mM/L represents the theoretical ionic strength of pure water, potential measurements are typically conducted in the presence of electrolytes and hence at higher ionic strengths.
Figure 3A shows that the surface charge density of -0.0007 e-/nm2 found from ALMO-EDA, which defines a true quantum mechanical definition of CT, would predict an unstable emulsion at any value of ionic strength, as the surface potential would be below --30 mV. Even under the assumption of a high surface charge density of -0.015 e-/nm2 the emulsions would be unstable at moderately high pH or salt strength, in contradiction to their experiments (i.e. oil droplet in 1mM NaOH or NaCl bulk solution)18. Oil-water emulsions are in fact stable over weeks to months8 at higher pH values, so CT cannot be the origin of oil charging as we have shown it is negligibly small.
[…]
Figure 3: Estimates of surface charge and zeta potentials. (A) Surface potential as a function of surface charge density ($\sigma_{0}$) for a 100 nm-radius spherical interface at various ionic strengths (1 mM, 0.01 mM, 0.0001 mM) using the non-linear Poisson Boltzmann equation34, 35, 36. Dashed vertical lines are surface charge densities of -0.015 17, -0.002 14, and -0.0007 e-/nm2 from ALMO-EDA. See Supplementary Information. (B). An illustration of the potential at the surface as assumed by a CT mechanism, and due to the double layer or beyond under the assumption of ionic species such as hydroxide ions.Adapted with permission from41. Copyright 2025 American Chemical Society
Pullanchery et al.18 also reported a simulation in which aqueous neopentane moved in the presence of an electric field. However, we re-analyzed the uploaded trajectories from their publication, finding that the observed net drift primarily corresponded to the center-of-mass motion of the entire simulated system, but not neopentane itself. This is because the oil carries no significant net charge from CT, and it thus explains why there is a lack of net drift under an applied electric field in the AIMD simulations of neopentane in water (see Supplementary Figure S2 and discussion).18
3 Discussion and Conclusion
Quantifying CT is rigorously possible using computational quantum chemistry calculations, which provide direct access to the electron density and even via the approximate wavefunction. Our results from ALMO-EDA19, 20, which defines a true quantum mechanical definition of CT as electron relaxation from an optimally polarized state to a delocalized state, allows for both forward and backward CT contributions to lowering the system’s energy. In the water dimer, CT is predominantly unidirectional, with the hydrogen bond acceptor oxygen transferring charge to the donor hydrogen, while backward CT is almost negligible. In contrast, forward CT in the water–hexane complex occurs from the hydrogen bond acceptor oxygen to the donor hydrogen in hexane, but this is largely balanced by backward CT from hexane to the hydrogen of water. This bidirectional nature of CT weakens the directionality of hydrogen bonding, contributing to the more flexible and distorted hydrogen-bonding geometry observed at the water–oil interface, and most relevant here, yields a negligible surface charge value of --0.0007 e-/nm2 that accumulates from oil to water. As we have shown, such insignificant amounts of surface charge would result in an instability of the oil-water emulsion, in contradiction to experimental observations.
Atomic charge partitioning methods assign electron density to individual atoms, which therefore includes a larger effect of partitioning the charge distribution of the optimally polarized state.19 This leads to an overestimation of the CT value and an unphysical estimate of the effective donor-acceptor energy gap42, 43. Hence the very large surface charge of --0.015 e-/nm2 from charge partitioning schemes used by Hassanali and co-workers are inherently flawed.17, 15 This is evident from the slow and smooth 1/r decay of charge accumulation deep into the oil phase in their theoretical analysis (see Figure 217 and Figure 818), behavior that instead signifies electrostatic screening rather than what is expected for true CT43 which is highly localized at the interface as we have shown above.
Since there is no net drift of the neopentane under an applied electric field, the electric-field-induced time-dependent polarization (also referred to as “dynamic polarization”) of neopentane cannot result in a sustained directional force as reported.18 Although dynamic polarization may give rise to transient forces when a molecule is exposed to an electric field, these effects are inherently short-lived, whereas the electrophoretic mobility reflects a steady-state phenomenon governed by the linear response of charges within the slip plane that defines the zeta potential. Regarding the pH dependence, it has been suggested that pH modifies electronic polarization of the bulk solution which induces mobility differences between hydroxide and chloride ions. However, it is known that Br- and I- both have greater polarizability than Cl- but there is almost no change in the mobility and zeta potential among the ions. A second explanation invokes the Grotthuss mechanism, suggesting that the conductivity of OH- is higher than that of Cl-. Yet the molar conductivity of OH- is lower than that of H3O+44; if mobility scales with conductivity, acidic solutions should exhibit higher mobility not hydroxide.
Furthermore, previous and unambiguous EDA studies have explicitly stated that CT would red shift, not blue shift, the C-H vibrational modes of hydrophobic groups interacting with water.25, 26, 22 Given the experimental observation by Roke and co-workers that their C-H vibrations blue shift, then CT as a mechanism must be ruled out, and instead is attributable to Pauli repulsion that arises as a confinement effect at the interface between oil and water. This is supported by the observation that the ”frozen” term in the ALMO-EDA is a majority of the energetic stabilization from C-H···O hydrogen bonds as we have shown in Figures 1 and 2. This is one of the necessary conditions for the VSFS spectroscopic features to be observed to blue-shift, i.e. relatively weak CT that is insufficient to compensate for the blue-shifting effect of the frozen interaction.22
Since the CT mechanism is clearly unviable as an explanation of emulsion stability and electrophoresis, as well as the VSFS measurements themselves, what is the molecular explanation for oil charge? The net oil charge must be large as it has been shown to give rise to large electric fields of $\sim$ --55 to --90 MV/cm at the interface.8 Assertions that impurities are responsible are also unlikely given the orders of magnitude increase in surface area of an emulsion compared to a planar interface. Furthermore their concentration would need to be large enough to exhibit some charging effect and yet at the same time such concentrations should be visible in the SFG and SCRS experiments, which they are not.15, 8
[…]
The implication of these related studies is that the surface potential and zeta potential are in fact not equivalent as assumed by Roke and co-workers, that ionic species are present and must be accounted for in the interpretation of their VSFS experiments, and clearly they should not be ignored in any physical models or simulations. At the same time the surface affinities of hydroxide and hydronium to simplified planar interfaces using simulation models are way too small to explain the needed surface charge densities to induce mobility and maintain emulsion stability. Hence it is still an open question at present as to the molecular origin that enhances the surface affinity of (presumably) hydroxide, but which no one has yet identified.
4 Methods
Charge Transfer Decomposition Analysis. The ALMO-EDA20 scheme separates the overall intermolecular interaction energy, $\Delta E_{\text{INT}}$, into contributions from frozen interaction ($\Delta E_{\text{FRZ}}$), polarization ($\Delta E_{\text{POL}}$), and charge transfer ($\Delta E_{\text{CT}}$) from the corresponding intermediate states
$\Delta E_{\text{INT}}=\Delta E_{\text{FRZ}}+\Delta E_{\text{POL}}+\Delta E_{\text{CT}}$ (3)
The frozen state is obtained by directly antisymmetrizing the isolated fragment wavefunctions. The polarization state is defined by the relaxation of the molecular orbitals of each fragment in the presence of other fragments and is CT-free.55, 56 In the large basis set limit, orbital overlap between fragments can cause charge transfer contamination in the polarized state. The second generation of ALMO-EDA resolved this by retaining only fragment electric response functions57, 58(FERFs), ensuring accurate polarized states even at the complete basis set limit. Finally, the CT state is achieved from the SCF calculation of the whole complex without any constraint, and it includes the electron flow to and from each molecule in the system due to intermolecular relaxation or mixing of the molecular orbitals.19, 29 The net direction of electron flow is determined by the balance between forward and backward contributions, processes that contribute to stabilizing the system by lowering its energy.
In the ALMO-EDA framework the amount of CT is defined as
$\Delta Q=\text{Tr}\{\hat{P}_{\text{CT}}\}-\text{Tr}\{\hat{P}_{\text{POL}}\hat{P}_{\text{CT}}\},$ (4)
where $\hat{P}_{\text{POL}}$ and $\hat{P}_{\text{CT}}$ are the density operators of the POL and CT states. $\Delta Q$ is the electron count in the final state that lies outside the density operator of the initial state, and is identical59 to the excitation number.60 The CT decomposition analysis of ALMO-EDA breaks this quantity into fragment additive pairs. For instance, the amount of charge transfer from fragment $x$ to fragment $y$ is obtained as
$\Delta Q_{x\rightarrow y}=\sum_{ia}\langle\psi^{ya}|\hat{P}^{\text{eff}}_{vo}|\psi_{xi}\rangle\langle\psi^{xi}|\hat{X}_{ov}|\psi_{ya}\rangle,$ (5)
where $\hat{P}^{\text{eff}}_{vo}$ is an effective density operator, $\hat{X}_{ov}$ is the generator of the unitary transformation connecting $\hat{P}_{\text{POL}}$ and $\hat{P}_{\text{CT}}$, and the $\psi$’s are the covariant and contravariant ALMOs. We can rotate the occupied and virtual ALMOs such that $\hat{X}_{ov}$ is diagonal under the rotated orbitals. These special orbitals form a most compact basis for describing the charge transfer, and they are called the complementary occupied-virtual pairs (COVP). Typically, the CT between two fragments in the CT process can be described by one or two dominant COVPs.
Given the balance between computational cost and accuracy, we selected def2-TZVPD as the optimal basis set for EDA (Details in Supplementary Figure S3). We note on technical grounds the basis set used by Poli and co-workers did not include diffuse functions17, although diffuse functions are necessary for properly describing electron density delocalization and hydrogen bond interactions, and their absence leads to misrepresenting the extent of CT in hydrogen-bonded networks61.
Due to the substantial size of the interfacial hexane–water clusters, which contain on average $\sim$400 atoms per cluster, performing EDA at the $\omega$B97X-V/def2-TZVPD level of theory poses a significant computational challenge, particularly for the evaluation of the polarized state. To alleviate the computational cost associated with polarization, we employed two key strategies to enhance efficiency. First, each hexane molecule was assigned as an independent fragment, thereby avoiding the overhead introduced by large fragments. All water molecules were treated collectively as a single fragment. Second, we utilized the recently developed uncoupled-FERFs (uFERFs)58, which provides an enormous speedup through a diagonal approximation of the orbital hessian matrix. The previous benchmark results 58 demonstrated that the polarization energies of neutral systems are almost unaffected by uFERF approximation.
Molecular dynamics simulations The system, consisting of 1450 water molecules and 200 hexane molecules, was simulated using the AMOEBA polarizable force field within the OpenMM62 platform and all calculations were done in double precision. The dimensions of the simulation box were set to 3.5 × 3.5 × 7.5 nm3. Simulations were kept at a constant temperature of 298.15 K and a constant Pzz of 1 atm, where Pzz is the component of the pressure tensor perpendicular to the interface (only the z-dimension of the periodic simulation box is allowed to change while the x- and y-dimensions are kept fixed). The simulations used the MTSLangevinIntegrator for efficient temperature control and integration. Non-bonded interactions were handled using the Particle Mesh Ewald (PME) method, with a non-bonded cutoff of 1.2 nm and an Ewald error tolerance of 0.0005. The system was equilibrated for 1 ns, with a time step of 0.5 fs. Following equilibration, a 20 ns production run was conducted to investigate the structural properties of the water-hexane interface. Snapshots were saved every 100 ps, yielding a total of 200 frames for EDA analysis.
[…]
DATA AVAILABILITY
AUTHOR CONTRIBUTIONS. R.Z. and T.H.-G. conceived the scientific direction of the project. R.Z. and H.S. performed all calculations. R.Z., H.S., M.H.-G., and T.H.-G performed analyses. R.Z. and T.H.-G. wrote the manuscript, and all authors provided comments on the results and manuscript.
ACKNOWLEDGMENT. We thank the CPIMS program, Office of Basic Energy Sciences, Chemical Sciences Division of the U.S. Department of Energy under Contract DE-AC02-05CH11231. MHG and HS acknowledge support from the U.S. National Science Foundation through Grant CHE-2313791. This work used computational resources provided by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract DE-AC02-05CH11231, and the Lawrencium computational cluster resource provided by the IT Division at the Lawrence Berkeley National Laboratory (Supported by the Director, Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231).